OpenStudy (anonymous):

what is the scientific notation of (6.7x10to the power of 4)(2.9x10to the power of 5)

OpenStudy (acxbox22):

(6.7 x 2.9)(10^4 x 10^5)

OpenStudy (mathstudent55):

\((6.7 \times 10^4) \times (2.9 \times 10^5) \) First, step, what is 6.7 * 2.9 = ?

OpenStudy (anonymous):

it would equal 19.43

OpenStudy (anonymous):

right?

OpenStudy (mathstudent55):

Good. Now what is \(10^4 \times 10^5\) = ?

OpenStudy (acxbox22):

yes and using exponent rules 10^4 x 10^5=10^9 put it all together and you get...???

OpenStudy (anonymous):

it would be 110,000 @mathstudent55

OpenStudy (anonymous):

and @acxbox22

OpenStudy (acxbox22):

no by put it together i mean 19.43 multiplied by 10^9 that is still not in scientific form so it is 1.943 x 10^9

OpenStudy (anonymous):

oh okay then hold up @acxbox22

OpenStudy (anonymous):

then it would be 110,019.43

OpenStudy (anonymous):

@acxbox22 you confusing me sorry i need step by step the way @mathstudent55 is doing

OpenStudy (mathstudent55):

No. There are two parts. 1. 6.7 * 2.9 = 19.43. You already did that. 2. 10^4 * 10^5 = 10^9 Now you put the two parts together, and you get this: That means \((6.7 \times 10^4) \times (2.9 \times 10^5) = 19.43 \times 10^9\) We're not done yet, though, because 19.43 is not scientific notation. The first part of the number in scientific notation has to be greater than or equal to 1 and less than 10. 19.43 is obviously greater than 10.

OpenStudy (anonymous):

i can see that 19.43 is greater than 10....

OpenStudy (mathstudent55):

Since to make 19.43 between 1 and 10 we need to divide 19.43 by 10, we need to multiply the second part, 10^9 by 10, to get 10^10. The answer is \(1.943 \times 10^{10}\)

OpenStudy (acxbox22):

whoops...it is 1.943 x 10^10 like @mathstudent55 said

OpenStudy (anonymous):

ohhh okay well thank you @mathstudent55

OpenStudy (anonymous):

do you guys want to help me with more or no?

OpenStudy (mathstudent55):

Sure. Pls start a new post.

OpenStudy (anonymous):

alright